OpenStudy (anonymous):

How many solutions does the following system have? y = -2/3 x + 3 and 2x + 3y = 9

OpenStudy (zzr0ck3r):

get them in the same form (2/3)x+y=3 2x+3y=9 lets get rid of the fraction on the first equation by multiplying each term by 3 we now have 2x+3y=9 and our second equation is 2x+3y=9 they are the exact same thing so they are the same line, so there are infinite solutions because there are infinite solutions to a line, and what ever satisfies the first line of course satisfies the second line

OpenStudy (anonymous):

Alright that makes sense, thanks!